OpenStudy (m_m13):

HELP!!!!!! will medal Which equations in point-slope form are equations of the line that pass through the points (4, 5) and (−3, −1)? Exactly two answers that are correct.

OpenStudy (m_m13):

A. y - 3 = 7/6 (x - 1) B. y - 5 = 6/7 (x - 4) C. y + 1 = 6/7 (x + 3) D. y - 4 = 7/6 (x - 5)

OpenStudy (m_m13):

@mathmate

OpenStudy (mathmate):

Point slope form: y-y1=m(x-x1) in this case, (x1,y1)=(4,5), (x2,y2)=(-3,-1) slope m = (y2-y1)/(x2-x1) Substitute the coordinates and you can find the choice that fits.

OpenStudy (nathalie0918):

it is equations b. and c.